# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

### v1

@@ Abstract.
Generative language modelling has surged in popularity with the emergence of services such as ChatGPT and Google Gemini. While these models have demonstrated transformative potential in productivity and communication, they overwhelmingly cater to high-resource languages like English. This has amplified concerns over linguistic inequality in natural language processing (NLP). This paper presents the first systematic review focused specifically on strategies to address data scarcity in generative language modelling for low-resource languages (LRL). Drawing from 54 studies, we identify, categorise and evaluate technical approaches, including monolingual data augmentation, back-translation, multilingual training, and prompt engineering, across generative tasks. We also analyse trends in architecture choices, language family representation, and evaluation methods. Our findings highlight a strong reliance on transformer-based models, a concentration on a small subset of LRLs, and a lack of consistent evaluation across studies. We conclude with recommendations for extending these methods to a wider range of LRLs and outline open challenges in building equitable generative language systems. Ultimately, this review aims to support researchers and developers in building inclusive AI tools for underrepresented languages, a necessary step toward empowering LRL speakers and the preservation of linguistic diversity in a world increasingly shaped by large-scale language technologies.
@@ 1. Introduction
Language modelling is a central task in the field of natural language processing (NLP) that aims to capture the features of a language by extracting the statistical and contextual relationships within a language (Wang et al., 2023). This information can subsequently be used to inform next-word predictions for a given sequence of previous terms (Kłosowski, 2018). It involves enlisting a model to extract the characteristics of a language from a large corpus of textual data that can consequently be used to process, understand and generate text (Thirunavukarasu et al., 2023). Language models evolved from statistical methods like Markov processes (Almutiri and Nadeem, 2022) and n-gram models (Roark et al., 2007) to recurrent neural network (RNN) approaches, such as Long Short-Term Memory (LSTM) RNNs (Van Houdt et al., 2020). However, the current rise in popularity of language models and language model-powered services such as ChatGPT (Lian et al., 2024; Ng and Chow, 2024) can be attributed to the breakthrough that introduced the transformer (Vaswani, 2017) as an architecture for modelling languages. This advancement enabled the development of pre-trained models such as the transformative Generative Pretrained Transformer (GPT) (Radford, 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin, 2018), with the transformer architecture enhancing NLP tasks such as machine translation (Raganato and Tiedemann, 2018), classification (McGiff and Nikolov, 2024) and named entity recognition (Arkhipov et al., 2019).
@@
A large volume of data is required to build a robust generative language model (Kaplan et al., 2020; Muennighoff et al., 2023; Hoffmann et al., [n. d.]; Radford et al., 2019). This conflicts with the limited textual data that is available to LRLs. Many studies aim to tackle the data disparity, with one main branch of these approaches focusing on integrating a data augmentation process into the model training pipeline (Shorten et al., 2021). Data augmentation is a process that generates synthetic data from an existing dataset (Shorten and Khoshgoftaar, 2019). The concept of data augmentation is commonly used in the field of computer vision to train more robust models (Taylor and Nitschke, 2018; Mikołajczyk and Grochowski, 2018; Khosla and Saini, 2020). Image datasets are often enlarged by performing image transformations such as rotations or brightness modifications to create additional training images (Mumuni and Mumuni, 2022). While it has been argued that NLP has not adopted data augmentation to the same degree as computer vision, a wide range of studies have enlisted a variety of processes to bridge the gap in available training data for LRL modelling (Shorten et al., 2021). Although prior surveys have explored data augmentation in NLP broadly, no systematic review to date has focused on how these strategies are applied in the development of generative models for LRLs.
@@ 2. Review Methodology
A systematic review is described by Page et al. as being a review that enlists clear, methodical techniques to aggregate and summarise the results of studies aimed at answering a clearly defined research question (Page et al., 2021). It is an informative process that can yield in-depth exploration and determine the current state of existing knowledge in a field. The PRISMA 2020 statement indicates that a systematic review can combine information relating to future research opportunities, answer research questions that span different individual studies, identify shortcomings in research and explore existing primary research results (Page et al., 2021). The PRISMA review methodology can be divided into four steps for this review: (1) Defining the research objectives/questions and eligibility criteria; (2) Adopting a search strategy; (3) Performing the study selection; and (4) Data extraction.
@@ 2.1. Defining the Research Questions and Eligibility Criteria
    The paper enlists a technical method for overcoming data scarcity.
@@ 3.2. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
This review has identified a wide range of LRLs that have been modelled by other studies. A total of 295 unique LRLs have been trained spanning a variety of language generation tasks and different model architectures. Figure 5 showcases the most frequently modelled LRLs by studies examined in this review. Bengali is the most frequently modelled LRL in this review, with 13% of papers including it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Shafayat et al., 2024; Bhowmick et al., 2023; Sorokin et al., 2022; Agarwal et al., 2022). Hindi follows behind Bengali with the language appearing in 11% of the studies (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Gerz et al., 2018; Maimaiti et al., 2021). Whereas, Tamil (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Gerz et al., 2018; Nissanka et al., 2020), Telugu (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Sorokin et al., 2022; Tanwar and Majumder, 2020; Dandapat and Federmann, 2018) and Turkish (Chang et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023) are each modelled by 9% of the papers, thus making them the third most commonly modelled LRLs. Although Turkish has been considered to be a high-resource language (Chang et al., 2023), some papers argue it is an LRL (Toraman, 2024; Yirmibeşoğlu and Güngör, 2023). Therefore we classify Turkish as an LRL in this review. Moreover, Javanese (Chang et al., 2023; Gerz et al., 2018; Cahyawijaya et al., 2021; Wongso et al., 2023) and Khmer (Agarwal et al., 2022; Gerz et al., 2018; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) have both appeared in 7% of the papers, with Azerbaijani (Chang et al., 2023; Pham et al., 2021; Maimaiti et al., 2021), Belarusian (Chang et al., 2023; Mao and Yu, 2024; Pham et al., 2021) and Slovak (Pham et al., 2021; Chang et al., 2023; Gerz et al., 2018) among the 16 languages that were modelled in 6% of studies times. Figure 5 is a word cloud that displays more of the most frequently modelled LRLs in this review. This word cloud highlights that LRLs such as Marathi (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Saini and Vidhyarthi, 2023) and Odia (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Agarwal et al., 2023) were also modelled somewhat frequently. Figure 6 reveals aggregations whereby each language is assigned to a category according to its respective language family. We have identified languages from 39 different language families that have been modelled by studies in this review. LRLs belonging to the Indo-European language family consist of 35% of the most frequently modelled LRLs. Austronesian languages are the second most commonly modelled LRL with a frequency of 15% in this review. Turkic and Niger-Congo-based languages are similarly represented with 8% and 7% frequency respectively. Uralic, Afro-Asiatic and Sino-Tibetan each represent 4% of the LRL families modelled by the literature in this review. Six remaining language families are represented with a frequency greater than 1%. These language families include Creole, Atlantic-Congo, Austro-Asiatic and Nakh-Daghestanian at 2%, with Mongolic and Constructed at 1%.
@@ 3.3.1. Monolingual Text Data Augmentation
The application of various strategies for bolstering monolingual text data occurs in 26% of the studies in this review. Such methods include generating new sentences for neural machine translation (NMT) by substituting low-frequency words with terms selected by an LSTM language model (Haque et al., 2021; Shi et al., 2022); constructing new sentences through the grammatical reordering of monolingual data using a reverse transcription grammar model (Wu and Wang, 2023); using a generative model combined with a grammar correction model to create grammatically correct synthetic data from the monolingual source data (Sun et al., 2019); and duplicating original sentences in a monolingual corpus and using a model to paraphrase them (Shi et al., 2022; Guo et al., 2021a; Gao et al., 2020). Other approaches include synthesising user-generated text for NMT by applying noise and stylistic changes to textual data (Marie and Fujita, 2020); appending document-specific metadata and information to the document itself (Usui and Komiya, 2023); generating text data for NMT by turning the process of augmentation into an optimisation problem (SwitchOut) (Haque et al., 2021; Shi et al., 2022; Wang et al., 2018); and using POS tagging to make replacements based on the label of the term (Maimaiti et al., 2021; Shi et al., 2022). Additional strategies involve boosting training data by reordering monolingual target language text to align with the structure of another language (Haque et al., 2021); incorporating knowledge at the word, phrase, and sentence levels to maximise the information gain from a monolingual corpus (Mi et al., 2024); and using soft contextual data augmentation to randomly replace a word with a similar meaning (Haque et al., 2021; Gao et al., 2019).
@@ 3.3.2. Back-Translation
Back-translation is also frequently used as a technique for addressing data scarcity, appearing in 24% of the studies in this review (Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Pham et al., 2021; Shi et al., 2022; Marie and Fujita, 2020; Wang et al., 2018; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Wang et al., 2022). This approach involves boosting monolingual data by using a backward NMT model to translate a monolingual corpus from a target language to a source language. This process of translating to another language and back again can produce variations that bolster the source language dataset. Although this technique starts with monolingual data and shares some characteristics with monolingual augmentation methods, this review treats back-translation as a distinct category due to its reliance on machine translation systems.
@@ 3.3.3. Prompt Engineering
A considerable portion of the studies in this review, specifically 20%, focused on addressing data scarcity by prompting existing generative language models. These existing models are examined based on their capacity to generate text for a wide range of LRLs. Several techniques have been proposed to improve generative language models. The programmer-interpreter approach iteratively refines a model’s output using a smaller language model for modifications (Li et al., 2024b). Another approach involves few-shot learning, where a model is taught unseen languages through examples and definitions (Rizqullah et al., 2023). Retrieval-augmented generation, which incorporates textbook-like material, enhances prompts and improves model output (Guo et al., 2024b). Some studies evaluate existing generative models on their effectiveness in handling low-resource languages (LRLs) such as Vallader (Bendel and Jabou, 2024), Odia (Agarwal et al., 2023), Faroese (Scalvini and Debess, 2024), and Marathi (Saini and Vidhyarthi, 2023). Instruction tuning is another category within our review paper set that tackles the data disparity problem by employing a technical approach. This method involves providing clear instructions for language models that enable them to adapt to unseen tasks and LRLs. In addition to adding machine translation instructions, word alignments have also been used to enhance the fine-tuning process (Mao and Yu, 2024). The efficacy of instruction tuning has also been assessed for adapting generative large language models for LRLs (Toraman, 2024). We have determined that two papers enlist cross-lingual information sharing as a method to bolster the data available for LRL models. Both papers build a cross-lingual information system that is responsible for collecting information for languages with more available training data when the information does not exist in the desired target language (Sorokin et al., 2022; Agarwal et al., 2022).
@@ 3.3.4. Multilingual Models
The category for multilingual models encapsulates all papers in this review that reference creating generative language models for more than a single language. 20% of the papers in this review were categorised for their use of multilingual models to address the issue of data scarcity when building generative models for LRLs. A number of NMT models naturally build generative language models that generate translations for pairs of languages (Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Guo et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Liu, 2022; Yu et al., 2020; Phan-Vu et al., 2017). In addition, several papers focused on building generative language models for a broader range of related languages. Some examples include a generative language model for 10 Indic languages (Niyogi and Bhattacharya, 2024); a model trained on a variety of Uralic languages (Downey et al., 2024); a multilingual model trained on a mixture of 250 high-resource and low-resource languages (Chang et al., 2023); and IndoNLG, which serves as a benchmark for generative tasks involving Indonesian languages (Cahyawijaya et al., 2021). Other notable multilingual models include a model paired with textbook-based prompts to teach LRLs to an existing language model (Guo et al., 2024b); a neural approach for next-word prediction across languages (Gerz et al., 2018); and a model that leverages cross-lingual information to extract content from high-resource languages like English and present it in another language (Sorokin et al., 2022; Agarwal et al., 2022). Additional work includes a translation model trained on a family of related Indonesian languages (Wongso et al., 2023), as well as a pre-trained language model designed for Chinese minority languages, such as Mongolian (Li et al., 2022a).

### v2

@@ Abstract.
Generative language modelling has surged in popularity with the emergence of services such as ChatGPT and Google Gemini. While these models have demonstrated transformative potential in productivity and communication, they overwhelmingly cater to high-resource languages like English. This has amplified concerns over linguistic inequality in natural language processing (NLP). This paper presents the first systematic review focused specifically on strategies to address data scarcity in generative language modelling for low-resource languages (LRL). Drawing from 54 studies, we identify, categorise and evaluate technical approaches, including monolingual data augmentation, back-translation, multilingual training and prompt engineering, across generative tasks. We also analyse trends in architecture choices, language family representation, and evaluation methods. Our findings highlight a strong reliance on transformer-based models, a concentration on a small subset of LRLs, and a lack of consistent evaluation across studies. We conclude with recommendations for extending these methods to a wider range of LRLs and outline open challenges in building equitable generative language systems. Ultimately, this review aims to support researchers and developers in building inclusive AI tools for underrepresented languages, a necessary step toward empowering LRL speakers and the preservation of linguistic diversity in a world increasingly shaped by large-scale language technologies.
@@ 1. Introduction
Language modelling is a central task in the field of natural language processing (NLP) that aims to capture the features of a language by extracting the statistical and contextual relationships within a language (Wang et al., 2023). This information can subsequently be used to inform next-word predictions for a given sequence of previous terms (Kłosowski, 2018). It involves using a model to extract the characteristics of a language from a large corpus of textual data that can consequently be used to process, understand and generate text (Thirunavukarasu et al., 2023). Language models evolved from statistical methods like Markov processes (Almutiri and Nadeem, 2022) and n-gram models (Roark et al., 2007) to recurrent neural network (RNN) approaches, such as Long Short-Term Memory (LSTM) RNNs (Van Houdt et al., 2020). However, the current rise in popularity of language models and language model-powered services such as ChatGPT (Lian et al., 2024; Ng and Chow, 2024) can be attributed to the breakthrough that introduced the transformer (Vaswani, 2017) as an architecture for modelling languages. This advancement enabled the development of pre-trained models such as the transformative Generative Pre-trained Transformer (GPT) (Radford, 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin, 2018), with the transformer architecture enhancing NLP tasks such as machine translation (Raganato and Tiedemann, 2018), classification (McGiff and Nikolov, 2024) and named entity recognition (Arkhipov et al., 2019).
@@
A large volume of data is required to build a robust generative language model (Kaplan et al., 2020; Muennighoff et al., 2023; Hoffmann et al., [n. d.]; Radford et al., 2019). This conflicts with the limited textual data that is available to LRLs. Many studies aim to tackle the data disparity, with one main branch of these approaches focusing on integrating a data augmentation process into the model training pipeline (Shorten et al., 2021). Data augmentation is a process that generates synthetic data from an existing dataset (Shorten and Khoshgoftaar, 2019). The concept of data augmentation is commonly used in the field of computer vision to train more robust models (Taylor and Nitschke, 2018; Mikołajczyk and Grochowski, 2018; Khosla and Saini, 2020). Image datasets are often enlarged by performing image transformations such as rotations or brightness modifications to create additional training images (Mumuni and Mumuni, 2022). While it has been argued that NLP has not adopted data augmentation to the same degree as computer vision, a wide range of studies have adopted a variety of processes to bridge the gap in available training data for LRL modelling (Shorten et al., 2021). Although prior surveys have explored data augmentation in NLP broadly, no systematic review to date has focused on how these strategies are applied in the development of generative models for LRLs.
@@ 2. Review Methodology
A systematic review is described by Page et al. as being a review that uses clear, methodical techniques to aggregate and summarise the results of studies aimed at answering a clearly defined research question (Page et al., 2021). It is an informative process that can yield in-depth exploration and determine the current state of existing knowledge in a field. The PRISMA 2020 statement indicates that a systematic review can combine information relating to future research opportunities, answer research questions that span different individual studies, identify shortcomings in research and explore existing primary research results (Page et al., 2021). The PRISMA review methodology can be divided into four steps for this review: (1) Defining the research objectives/questions and eligibility criteria; (2) Adopting a search strategy; (3) Performing the study selection; and (4) Data extraction.
@@ 2.1. Defining the Research Questions and Eligibility Criteria
    The paper uses a technical method for overcoming data scarcity.
@@ 3.2. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
This review has identified a wide range of LRLs that have been modelled by other studies. A total of 295 unique LRLs have been trained spanning a variety of language generation tasks and different model architectures. Figure 5 showcases the most frequently modelled LRLs by studies examined in this review. Bengali is the most frequently modelled LRL, with 13% of papers including it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Shafayat et al., 2024; Bhowmick et al., 2023; Sorokin et al., 2022; Agarwal et al., 2022). Hindi follows behind Bengali with the language appearing in 11% of the studies (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Gerz et al., 2018; Maimaiti et al., 2021). Whereas, Tamil (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Gerz et al., 2018; Nissanka et al., 2020), Telugu (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Sorokin et al., 2022; Tanwar and Majumder, 2020; Dandapat and Federmann, 2018) and Turkish (Chang et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023) are each modelled by 9% of the papers, thus making them the third most commonly modelled LRLs. Although Turkish has been considered to be a high-resource language (Chang et al., 2023), some papers argue it is an LRL (Toraman, 2024; Yirmibeşoğlu and Güngör, 2023). Therefore we classify Turkish as an LRL. Moreover, Javanese (Chang et al., 2023; Gerz et al., 2018; Cahyawijaya et al., 2021; Wongso et al., 2023) and Khmer (Agarwal et al., 2022; Gerz et al., 2018; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) have both appeared in 7% of the papers, with Azerbaijani (Chang et al., 2023; Pham et al., 2021; Maimaiti et al., 2021), Belarusian (Chang et al., 2023; Mao and Yu, 2024; Pham et al., 2021) and Slovak (Pham et al., 2021; Chang et al., 2023; Gerz et al., 2018) among the 16 languages that were modelled in 6% of studies times. Figure 5 is a word cloud that displays more of the most frequently modelled LRLs. This word cloud highlights that LRLs such as Marathi (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Saini and Vidhyarthi, 2023) and Odia (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Agarwal et al., 2023) were also modelled somewhat frequently. Figure 6 reveals aggregations whereby each language is assigned to a category according to its respective language family. We have identified languages from 39 different language families that have been modelled by studies in this review. LRLs belonging to the Indo-European language family consist of 35% of the most frequently modelled LRLs. Austronesian languages are the second most commonly modelled LRL with a frequency of 15%. Turkic and Niger-Congo-based languages are similarly represented with 8% and 7% frequency respectively. Uralic, Afro-Asiatic and Sino-Tibetan each represent 4% of the LRL families modelled. Six remaining language families are represented with a frequency greater than 1%. These language families include Creole, Atlantic-Congo, Austro-Asiatic and Nakh-Daghestanian at 2%, with Mongolic and Constructed at 1%.
@@ 3.3.1. Monolingual Text Data Augmentation
The application of various strategies for bolstering monolingual text data occurs in 26% of the studies. Such methods include generating new sentences for neural machine translation (NMT) by substituting low-frequency words with terms selected by an LSTM language model (Haque et al., 2021; Shi et al., 2022); constructing new sentences through the grammatical reordering of monolingual data using a reverse transcription grammar model (Wu and Wang, 2023); using a generative model combined with a grammar correction model to create grammatically correct synthetic data from the monolingual source data (Sun et al., 2019); and duplicating original sentences in a monolingual corpus and using a model to paraphrase them (Shi et al., 2022; Guo et al., 2021a; Gao et al., 2020). Other approaches include synthesising user-generated text for NMT by applying noise and stylistic changes to textual data (Marie and Fujita, 2020); appending document-specific metadata and information to the document itself (Usui and Komiya, 2023); generating text data for NMT by turning the process of augmentation into an optimisation problem (SwitchOut) (Haque et al., 2021; Shi et al., 2022; Wang et al., 2018); and using POS tagging to make replacements based on the label of the term (Maimaiti et al., 2021; Shi et al., 2022). Additional strategies involve boosting training data by reordering monolingual target language text to align with the structure of another language (Haque et al., 2021); incorporating knowledge at the word, phrase, and sentence levels to maximise the information gain from a monolingual corpus (Mi et al., 2024); and using soft contextual data augmentation to randomly replace a word with a similar meaning (Haque et al., 2021; Gao et al., 2019).
@@ 3.3.2. Back-Translation
Back-translation is also frequently used as a technique for addressing data scarcity, appearing in 24% of the studies (Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Pham et al., 2021; Shi et al., 2022; Marie and Fujita, 2020; Wang et al., 2018; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Wang et al., 2022). This approach involves boosting monolingual data by using a backward NMT model to translate a monolingual corpus from a target language to a source language. This process of translating to another language and back again can produce variations that bolster the source language dataset. Although this technique starts with monolingual data and shares some characteristics with monolingual augmentation methods, this review treats back-translation as a distinct category due to its reliance on machine translation systems.
@@ 3.3.3. Prompt Engineering
A considerable portion of the studies, specifically 20%, focused on addressing data scarcity by prompting existing generative language models. These existing models are examined based on their capacity to generate text for a wide range of LRLs. Several techniques have been proposed to improve generative language models. The programmer-interpreter approach iteratively refines a model’s output using a smaller language model for modifications (Li et al., 2024b). Another approach involves few-shot learning, where a model is taught unseen languages through examples and definitions (Rizqullah et al., 2023). Retrieval-augmented generation, which incorporates textbook-like material, enhances prompts and improves model output (Guo et al., 2024b). Some studies evaluate existing generative models on their effectiveness in handling low-resource languages (LRLs) such as Vallader (Bendel and Jabou, 2024), Odia (Agarwal et al., 2023), Faroese (Scalvini and Debess, 2024), and Marathi (Saini and Vidhyarthi, 2023). Instruction tuning is another category within our review paper set that tackles the data disparity problem by employing a technical approach. This method involves providing clear instructions for language models that enable them to adapt to unseen tasks and LRLs. In addition to adding machine translation instructions, word alignments have also been used to enhance the fine-tuning process (Mao and Yu, 2024). The efficacy of instruction tuning has also been assessed for adapting generative large language models for LRLs (Toraman, 2024). We have determined that two papers use cross-lingual information sharing as a method to bolster the data available for LRL models. Both papers build a cross-lingual information system that is responsible for collecting information for languages with more available training data when the information does not exist in the desired target language (Sorokin et al., 2022; Agarwal et al., 2022).
@@ 3.3.4. Multilingual Models
The category for multilingual models encapsulates all papers that reference creating generative language models for more than a single language. 20% of the studies were categorised for their use of multilingual models to address the issue of data scarcity when building generative models for LRLs. Several NMT models naturally build generative language models that generate translations for pairs of languages (Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Guo et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Liu, 2022; Yu et al., 2020; Phan-Vu et al., 2017). In addition, several papers focused on building generative language models for a broader range of related languages. Some examples include a generative language model for 10 Indic languages (Niyogi and Bhattacharya, 2024); a model trained on a variety of Uralic languages (Downey et al., 2024); a multilingual model trained on a mixture of 250 high-resource and low-resource languages (Chang et al., 2023); and IndoNLG, which serves as a benchmark for generative tasks involving Indonesian languages (Cahyawijaya et al., 2021). Other notable multilingual models include a model paired with textbook-based prompts to teach LRLs to an existing language model (Guo et al., 2024b); a neural approach for next-word prediction across languages (Gerz et al., 2018); and a model that leverages cross-lingual information to extract content from high-resource languages like English and present it in another language (Sorokin et al., 2022; Agarwal et al., 2022). Additional work includes a translation model trained on a family of related Indonesian languages (Wongso et al., 2023), as well as a pre-trained language model designed for Chinese minority languages, such as Mongolian (Li et al., 2022a).